# Vikhr: Constructing a State-of-the-art Bilingual Open-Source Instruction-Following Large Language Model for Russian

### v4

@@ 3.3 Instruction Tuning
En-MMLU Ru-MMLU CheGeKa
@@
MTS AI Chat 7B (closed-source) $\diamondsuit$ false sft+dpo - 0.689 0.083 0.56 0.479
GigaChat-7B (closed-source) $\diamondsuit$ true sft+dpo - 0.67 0.451* 0.71* 0.479
aya101 false pt+sft 0.41 0.37 0.005 0.36 0.320
Mistral-7B-Instruct-v0.2 false none 0.60 0.78 0.005 0.57 0.400
rccmsu/ruadapt-mistral-7b-v0.1 false pt+sft 0.61 0.72 0.005 0.64 0.421
rugpt13b true none 0.25 0.25 0.132 0.52 0.208
saiga-mistral-7b-lora false sft 0.60 0.76 0.223 0.64 0.442
saiga-llama3-8b false sft 0.59 0.78 0.225 0.66 0.476
(Ours) Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.231 0.67 0.485
@@ 4.1 Experimental Setup
The evaluation was performed on MMLU (En-MMLU) Hendrycks et al. (2021), Ru-MMLU [^6] https://huggingface.co/datasets/NLPCoreTeam/mmlu_ru, Ru-MMLU-pro [^7] https://huggingface.co/datasets/Vikhrmodels/mmlupro-ru, ruXNLI Conneau et al. (2018), CheGeKa Mikhalkova and Khlyupin (2022), Russian SuperGLUE Shavrina et al. (2020), and MERA Fenogenova et al. (2024). The MMLU (En-MMLU) benchmark assesses LLMs across 57 subjects through multiple-choice questions, measuring a model’s general knowledge and reasoning abilities. We utilize this benchmark to verify that the model retains its bilingual proficiency. For this dataset, we report the accuracy@1 score. We constructed Ru-MMLU and Ru-MMLU-pro by automatically translating original English MMLU and MMLU-pro to Russian. Translation was done by GPT-3.5 and GPT-4 respectively. Just as for MMLU, for these datasets, we report the accuracy@1 score. CheGeKa is based on the questions from the Russian version of “Jeopardy” (the “Own game”). This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks, MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.

### v5

@@ 3.3 Instruction Tuning
En-MMLU Ru-MMLU
Ru-MMLU
pro
CheGeKa
@@
MTS AI Chat 7B (closed-source) $\diamondsuit$ false sft+dpo - 0.69 - 0.083 0.56 0.479
GigaChat-7B (closed-source) $\diamondsuit$ true sft+dpo - 0.67 - 0.451* 0.71* 0.479
aya101 false pt+sft 0.41 0.37 0.12 0.005 0.36 0.320
Mistral-7B-Instruct-v0.2 false none 0.60 0.78 0.11 0.005 0.57 0.400
rccmsu/ruadapt-mistral-7b-v0.1 false pt+sft 0.61 0.72 0.16 0.005 0.64 0.421
rugpt13b true none 0.25 0.25 0.10 0.132 0.52 0.208
saiga-mistral-7b-lora false sft 0.60 0.76 0.14 0.223 0.64 0.442
saiga-llama3-8b false sft 0.59 0.78 0.22 0.225 0.66 0.476
(Ours) Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.19 0.231 0.67 0.485
@@ 4.1 Experimental Setup
The evaluation was performed on MMLU (En-MMLU) Hendrycks et al. (2021), Ru-MMLU [^6] https://huggingface.co/datasets/NLPCoreTeam/mmlu_ru, Ru-MMLU-pro [^7] https://huggingface.co/datasets/Vikhrmodels/mmlupro-ru, ruXNLI Conneau et al. (2018), CheGeKa Mikhalkova and Khlyupin (2022), Russian SuperGLUE shavrina2020russiansuperglue, and MERA fenogenova2024mera. The MMLU (En-MMLU) benchmark assesses LLMs across 57 subjects through multiple-choice questions, measuring a model’s general knowledge and reasoning abilities. We utilize this benchmark to verify that the model retains its bilingual proficiency. For this dataset, we report the accuracy@1 score. We constructed Ru-MMLU and Ru-MMLU-pro by automatically translating original English MMLU and MMLU-pro to Russian. Translation was done by GPT-3.5 and GPT-4 respectively. Just as for MMLU, for these datasets, we report the accuracy@1 score. CheGeKa is based on the questions from the Russian version of “Jeopardy” (the “Own game”). This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks, MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.